Oxford Materials IT Committee Meeting

MICHAELMAS TERM 2013
A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 29th October 2013 in the PR Meeting Room.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters oustanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 4th February 2014 at 11am in PR meeting room

ITEM 2. MATTERS OUTSTANDING FROM THE PREVIOUS MEETING

See minutes below.

ITEM 3: MATTERS FOR DISCUSSION

XP end-of-life planning – status update

ITEM 4: HEAD OF IT REPORT

The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: Rob Saunders joined the IT support team replacing Alex Cock, although Alex is still providing 40% FTE remote support for the server infrastructure.

Admin Network Scientia timetabling system has been dropped again after only one term due to lack on on-going technical support. Debbie Townsend replaced Hettie Gonzalez.

Teaching Laboratory The 6 microscope control computers have been replaced and are now running Windows 7. The laptops have had their hard disks replaced and are also running Windows 7, but are still very old and slow and are still in need of replacement when funds permit. The computer room desktops have not yet been upgraded due to technical dependencies. We plan to replace/upgrade those desktops over the Christmas vacation. Matlab and Labview have been installed on the 6 microscope computers, as required for a new 1st year practical.

Lecture theatres / meeting rooms All rooms checked at end of term and before start of term. There have been fewer problems with whiteboards becoming dirty since Dawn Hazel has been doing weekly checks of all rooms. The video conference equipment in the Banbury Road lecture theatre is now permanently installed and can selected via the lectern projector control panel. A plan is being formulated for upgrading the AV equipment in the Parks Road meeting room.

Digital Print-room Numerous hardware problems with the colour printer have been resolved.

Library Nothing to report.

MML All centrally managed Opensuse desktops have migrated to being standalone Ubuntu.desktops. Eight new compute nodes were purchased to complete the cluster Ironman. Four new servers have been commissioned as code development systems.

EM A new UPS system was installed on the Jeol2200. Major rewiring of the basement of the Holder Building is almost complete. All networking has been replaced. Door entry system has been kept due to lack of funds to replace. 10 phones have been migrated to VOIP.

MFFP Synology storage array expanded by a further 40Tb (now full capacity). Banbury Road lecture theatre was modified to make the video conference equipment permanent and the DTN lectures are currently being broadcast without problem so far.

OMCSimplications of XP end-of-life upon all equipment control systems is still being reviewed.

Website

Security updates were applied to (11) cmsmadesimple sites twice. However www.materials and www-solar.materials still need attention in order to be upgradable to the current version of CMS made simple. Recently live websites: defects.materials (Fitzgerald group), oxfordpes2013.materials (Giustino conference), jcmaxwell.materials (Yates project), www-eia.materials (Kirkland group), atomprobe.materials (Moody group) Other sites still under development: www.materials (Marrow project), qnano.materials (GADB project), www-emtest.materials (EM instruments website), The 2days per week web support from Kay Davies which started in April petered out again by the end of May.

Publications 19 sets of business cards and several visa passport photographs. Departmental handbooks were printed (at short notice again!) ready for the new academic year.

Network New wireless access point and 12 new wired ports in Hume-Rothery basement. New network cabinet and building core in Holder Building plus full rewire of ground floor. Several network loops detected and resolved. Begbroke core switch was replaced due to switch memory failure. Begbroke network UPS batteries changed. New wireless access point installed in Hirsch North. Grobert group purchased 10 additional dual-network routers.

Security No significant infections or security breaches detected.

Disposal A large quantity of waste electrical equipment was disposed at the start of October: 1x UPS; 13x CRT; 8x flatscreens; 57x computers; 14x printers; 9x switches; 1x microwave oven; 11x boxes misc + keyboards etc.

Software Origin university-wide site license was renewed from 5th August 2013 for 3 years. Hardware instability of the rothery2 in September affected availability of some software licenses (Windows Fortran, Materials Studio, Mathematica8). Rothery3 now hosts licences for Mathematica9, COMSOL(RF module, ACDC modules). Scientia timetabling system was decommissioned again after only one term due to lack of on-going support. New versions of Matlab2013b and Abaqus6.13 are available after upgrade of the central license server. The central pool of Matlab licenses for the university has been expanded from 450 to 550 licenses but the cost of a full site license is still prohibitive. Annual Departmental contributions to site licenses for Mathematica, Matlab student version and Labview were paid.

Disaster Recovery and Backup There were serious problems with stability of the departmental file server at the start of August. Eventually the issues were identified and resolved but service availability dropped to 89% between 1-14th August compared to >99.9% availability normally. Jeol6500 disk failed but recovered from ghost backup.

New Device registrations May 16; June 12; July 16; August 7; September 57; October 29;


MINUTES

A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 7th May 2013 in the PR Meeting Room. The following people attended: APD, PJW, RS, KSc, CA, RT, SCB, PE, KSu, AL
Apologies: Jonathan Yates

1. Apologies for Absence and Membership of the Committee

APD welcomed Rob Saunders who started work on 13th May 2013 replacing Alex Cock.

Committee membership was reviewed and is up to date.

2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) IT disaster recovery plans (Minute 2a)
IT staff have implemented a private mediawiki installation to store IT information including notes about all tasks and configurations. The system automatically manages versions and dates and has a good search interface for retrieving information. A site structure has been created and information is gradually being populated as tasks are carried out. One day a month has been allocated towards documentation updating to include more details about servers and services including service/server location, failover design, backup location(s) and specific configurations. Disaster recovery procedures need to be documented.

b) *Podcasting (Minute 2b)
SCB has the video files for Sabina Caneva’s PartII talk last year but we need an mechanism for annually paying someone to generate a podcast video of the prize winning Part II and graduate presentations. PJW will arrange recording of this year’s winner Lucy Durrans via OUCS. (No progress)

Action: PJW

c) *PC maintenance – OUCS computer hardware breakdown service (Minute 2c)

A process is still required for (automatic?) notification of owners that their warranty is due to expire and directing them to the IT services hardware breakdown scheme if warranty is required – such as for laptops. (No progress)

Action: PJW

d) *IT strategic plan (Minute 2d)
PJW still needs to rework the draft strategy document into a questionnaire. (No progress)

Action: PJW/APD

e) Server Infrastructure upgrade (Minute 2e)

The problem with incomplete file uploads via secure shell has been resolved. A problem with file upload via Netstorage has been discovered but it seems to only affect Chrome web browser and is thus easily circumvented.

A pair of virtual host servers and a pair of network switches were purchased and installed and all production virtual machines have been migrated onto the new hardware. The old virtual host servers are available for redeployment, possibly as terminal servers. There was some discussion of the potential benefits of terminal servers. The old windows update server and old Sophos enterprise console server is almost ready to decommission (oums-wsus). After a recent push at migrating the last few systems there now only 3 desktops talking to Sophos and 11 desktops talking to WSUS. The old Altiris management server (oums-nomad) has been virtualized. AMC is now working on migrating the non-user data (e.g. applications) such that the old server can be safely turned off.

Action: AMC

f) *Photocopying and Printing (Minute 2g)
When time permits the network printers will be migrated from the old iPrint server with Pcounter print monitoring to the new iPrint servers with Papercut-MF print monitoring.

Action: AMC

g) Draft Information Security Policy (Minute 3b)
(No progress since previous meeting.)
DC appointed PJW as the Departmental Information Security Officer (DISO), responsible for managing Information Security within the Department of Materials. Other staff members specifically identified as being well placed to help PJW with the formulation of a Departmental Information Security Policy were: APD (administrative and personnel information), AOT (student information), JRY( college tutor information, research information ), JAAC (research information, experience with ISO certification standards). PJW presented a brief action plan for developing the required Department Information Security Policy. DC were keen that PJW should simplify the process for staff to supply details about their information assets (e.g. short questionnaire) and that PJW should provide clear guidance to staff about best practice information handling. Kay Davies to assist collating data.

Action: PJW

h) *Finances (Minute 5)
PJW to circulate financial breakdown for 2011-2012 to the committee. (No progress)

Action: PJW

i) *EM bookings (Minute 6b)
PJW EG reported that JCCG had requested the equipment booking system be modified to incorporate an instrument status traffic light scheme indicating whether the instrument was fully functional or not. PJW is already working on re-writing the booking system and will attempt to incorporate such a feature. The updated system is still not ready to go live.APD reported that Neil Young is trialling some commercial booking software.

Action: PJW

j) Teaching laboratory computer upgrades (Minute 3a)
Adrian Taylor presented the teaching laboratory review plans to DC in Trinity Term. Since that meeting the 6 microscope control computers have been replaced and are now running Windows 7. The laptops have had their hard disks replaced and are also running Windows 7, but are still very old and slow and are still in need of replacement when funds permit. The computer room desktops have not yet been upgraded due to technical dependencies. We plan to replace/upgrade those desktops over the Christmas vacation. Matlab and Labview have been installed on the 6 microscope computers, as required for a new 1st year practical.

3 Matters for Discussion

3a) XP end of life planning
PJW presented an updated discussion document planning for XP end-of-life in April 2014. IT staff have performed ~93 installations to Windows7 since the last meeting (both new and old equipment). A database query for the number of XP computers seen in the past 2 months now lists about 210 systems, down from 316. Many of these computers (~90) are Novell managed desktops which require infrastructure changes before being able to migrate to Windows7 (we hope to upgrade Zenworks by December). Also many of these computers are old hardware unable to upgrade to Windows 7 or equipment control systems unable to upgrade to Windows 7.

Therefore we have been making steady progress towards upgrading those computers able to run Windows7 but there is still a lot to do before April 2014.

APD full requested a list of computers still using Windows XP, in particular equipment computers, so that the risks can be reviewed and so that at the next meeting ITC can define a policy concerning what to do with XP computers that cannot be upgraded for whatever reason.

Action: PJW

3b) Office 2003 end of life planning
The number of computers running Office 2003 (in the past 2 months) has decreased from 106 to around 48. It should be relatively straight forward to update the remaining systems. IT staff are currently deploying Office 2013 on new installations.

RT noted that users upgrading Office should be directed to training courses provided by central IT services and links to online resources to aid the transition.

Action: PJW

3c) Novell managed desktops
RT queried our current dependency upon Novell. PJW noted that we have used Novell for provision of desktop computing to the teaching laboratory and to administrative staff since the mid 1990s, initially managed via NSMS and subsequently managed by Alex Cock. However the departure of Alex and arrival of Rob who has less experience with Novell provides an opportunity to plan to migrate away from Novell managed desktops. Rob is currently developing an alternative managed desktop environment based upon Microsoft technology. However since the timescale for being able to transition is still unclear the Zenworks upgrade necessary to support Novell with Windows7 is also being implemented by Alex.

Action: RS/AMC

4. Head of IT Report
A written report was circulated with the agenda.

5. Finances
No information was presented.

6. AOB
No matters arising.

7. Date of Next Meeting
The next meeting is on Tuesday 4th February 2014 at 11am in PR meeting room.
The meeting closed at 12.00pm.